### v1

# DiffX: Guide Your Layout to Cross-Modal Generative Modeling

Abstract
Diffusion models have made significant strides in text-driven and layout-driven image generation. However, most diffusion models are limited to visible RGB image generation. In fact, human perception of the world is enriched by diverse viewpoints, including chromatic contrast, thermal illumination, and depth information. In this paper, we introduce a novel diffusion model for general layout-guided cross-modal “RGB+X” generation, called DiffX. We firstly construct the cross-modal image datasets with text descriptions using the LLaVA model for image captioning, supplemented by manual corrections. Notably, DiffX presents a simple yet effective cross-modal generative modeling pipeline, which conducts diffusion and denoising processes in the modality-shared latent space, facilitated by our Dual-Path Variational AutoEncoder (DP-VAE). Furthermore, we incorporate the gated cross-attention mechanism to connect the layout and text conditions, leveraging Long-CLIP for embedding long captions to enhance user guidance. Through extensive experiments, DiffX demonstrates robustness and flexibility in cross-modal generation across three RGB+X datasets: FLIR, MFNet, and COME15K, guided by various layout types. It also shows the potential for adaptive generation of “RGB+X+Y” or more diverse modalities. Our code and processed image captions are available at https://github.com/zeyuwang-zju/DiffX.
1 Introduction
Over the past decade, there have been significant advancements in image generation, driven by the improvement of deep generative models such as Variational AutoEncoders (VAEs) (Kingma and Welling 2013) and Generative Adversarial Networks (GANs) (Goodfellow et al. 2014). Notably, the recent diffusion models (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020) like DALL-E (Ramesh et al. 2021, 2022), Imagen (Saharia et al. 2022), and Stable Diffusion (SD) (Rombach et al. 2022) have gained popularity for their ability to generate high-fidelity images conditioned on text descriptions or class labels. Moreover, the field has witnessed the rise of layout-to-image models based on various layout conditions, including bounding boxes, semantic maps, and keypoints (Xie et al. 2023; Li et al. 2023; Ma et al. 2024). The success of diffusion models can be attributed to its stable learning objective and the availability of large-scale multi-modal training data.
Figure 1: Cross-modal “RGB+X” generation in latent space.
Figure 2: (a) The process of preparing the image captions. (b) Examples of cross-modal images, labels, and prepared captions.
Current layout-to-image diffusion models primarily focus on generating high-quality visible RGB images. However, our perception of the world is enriched by various perspectives beyond the visible spectrum, such as thermal illumination and depth information. Cross-modal visual understanding typically leverages the input format of “RGB+X”, where X represents additional data like Thermal (T) or Depth (D) images. It offers a more comprehensive representation compared with RGB-only understanding, especially in challenging environments. At present, a significant limitation lies in the scarcity of training data containing pixel-aligned RGB+X image pairs. If the conventional generative models are adopted for cross-modal data augmentation, we can only generate RGB and X images separately, and as a result, the RGB and X images cannot form consistent pairs of matching data. Consequently, we wonder if we can generate the cross-modal images simultaneously under layout guidance by an integrated model?
In this work, we introduce a novel diffusion model for cross-modal image generation, called DiffX. As shown in Figure 1, the proposed DiffX is capable of generating RGB+X image pairs guided by various layouts with user instructions. Our main contributions are as follows:
  • Due to the limited cross-modal images with text descriptions, we leverage the Large-Multimodal Model (LMM) to prepare the captions, followed by manual corrections.
  • We propose an effective cross-modal generative modeling pipeline, which performs the diffusion and denoising processes in the modality-shared latent space. Specially, the DP-VAE is implemented for cross-modal encoding.
  • We employ the gated cross-attention mechanism to establish a comprehensive connection between layout and text conditions, where the Long-CLIP (Zhang et al. 2024) is adopted to capture critical information in captions.
  • Experiments demonstrate that DiffX is capable of generating high-quality and coherent RGB+X image pairs based on bounding boxes, semantic maps, salient maps, and edge contours. Additionally, it demonstrates adaptability to “RGB+X+Y” or diverse-modality generation.
Figure 3: The workflow of our DiffX model for cross-modal generation. It performs the diffusion and denoising processes in the modality-shared space. Finally, the denoised feature $\boldsymbol{z}_{0}$ is decoded into the “RGB+X” image pair via the dual-path decoders.
2 Related Works
2.1 Cross-Modal Visual Understanding
Cross-modal visual understanding normally adopts RGB+X input data. Thermal images have shown advantages in detecting objects in low-light conditions. For instance, CFT (Fang, Han, and Wang 2021) and ICAFusion (Shen et al. 2024) incorporate Transformer-based RGB+T feature fusion modules for multispectral object detection. In semantic segmentation, EGFNet (Zhou et al. 2022) leverages prior edge maps of RGB+T images to capture detailed contour information. Secondly, depth images offer valuable insights into the 3D layouts and spatial structures. For example, HidaNet (Wu et al. 2023) utilizes a granularity-based attention mechanism to enhance the advantages of RGB+D features for Salient Object Detection (SOD). However, current cross-modal visual understanding is hindered by the lack of well-aligned RGB+X image pairs for training deep models.
2.2 Layout-to-Image Generation
Based on Denoising Diffusion Probabilistic Model (DDPM) (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020), diffusion models have undergone advancements in training and sampling techniques (Song, Meng, and Ermon 2021; Nichol and Dhariwal 2021). Recent studies have proposed layout-to-image diffusion models, which provide precise instructions for the object placement. Notably, LAW-Diffusion (Yang et al. 2023) integrates a spatial parser and adaptive guidance, enabling complex scene generation. BoxDiff (Xie et al. 2023) directly incorporates spatial conditions into the training-free denoising process. InstanceDiffusion (Wang et al. 2024) focuses on precise instance-level control in image generation.
2.3 Multi-Modal Generation
Multi-modal generation has become a challenging task and attracted widespread research. MM-Diffusion (Ruan et al. 2023) integrates featuring two interconnected denoising autoencoders and a sequential multi-modal U-Net architecture for joint audio-video generation. Moreover, MM-Interleaved (Tian et al. 2024) dynamically extracts information from multi-scale features for image-text generation. In addition, ControlVAR (Li et al. 2024), MT-Diffusion (Chen et al. 2024), and 4M (Mizrahi et al. 2024) adopt large-scale natural datasets or private datasets for end-to-end training on conditional generation. Despite these advancements, there remains a notable gap in generative models focused on layout-guided cross-modal “RGB+X” image generation.
3 LMM-Assisted Caption Generation
For training our diffusion model for cross-modal generation, it is necessary to prepare detailed captions for images, where the workflow is illustrated in Figure 2 (a).
In detail, the first step involves the utilization of the advanced LMM, namely LLaVA-v1.5-7b model (Liu et al. 2024), to extract descriptive information related to weather conditions, environmental settings, surrounding elements, transportation conditions, and identifiable objects within the RGB images. Despite the effectiveness of LLaVA in capturing visual content, it is prone to inaccuracies and omissions captioning, especially in complex scenes. Additionally, it lacks the ability to simultaneously annotate complementary information in both RGB and X images. Therefore, a correction procedure is implemented after the generation of raw captions. In cases where the generated captions exhibit minimal errors, manual corrections are applied to enhance accuracy. However, instances of significant errors or insufficient descriptions prompting ambiguity necessitate further corrections. Under such conditions, the GPT-4 language model (Achiam et al. 2023) serves as a valuable tool for providing supplementary descriptions and rectifying inaccurate information to ensure a comprehensive image depiction. By combining the capability of the LLaVA model with the manual correction and assistance of GPT-4, the captioning process achieves a high level of fidelity and accuracy in conveying the visual context in the cross-modal datasets.
4 DiffX for Cross-Modal Generation
In this section, we introduce our proposed DiffX, designed to generate cross-modal “RGB+X” images, as illustrated in Figure 3. We first introduce the preliminaries on Latent Diffusion Model (LDM). Then, we describe the workflow and components of our DiffX in detail. Finally, we provide a discussion on the advantages and unique features of DiffX.
4.1 Preliminaries on Latent Diffusion Model
Building upon DDPM, LDM and the successor SD (Rombach et al. 2022) have pushed the boundaries of diffusion models by using a low-dimensional latent space. Inside LDM, a prior VAE is employed to extract the latent representation $\boldsymbol{z}$ of the ground-truth image $\boldsymbol{I}$. Subsequently, the diffusion model $\boldsymbol{\epsilon}_{\boldsymbol{\vartheta}}$, normally implemented as a UNet (Ronneberger, Fischer, and Brox 2015) with residual (He et al. 2016) and self-attention (Vaswani et al. 2017) blocks, is adopted to denoise the anormally-distributed $\boldsymbol{\epsilon}$ in the noisy latent $\boldsymbol{z}_{t}$ at time step $t$. The training objective can be represented as:
$\min_{\boldsymbol{\theta}}\mathcal{L}_{\text{LDM}}=\mathbb{E}_{\boldsymbol{z}_{t},\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t,\boldsymbol{c}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\boldsymbol{\vartheta}}\left(\boldsymbol{z}_{t},t,\boldsymbol{c}\right)\right\|_{2}^{2}\right],$ (1)
where $\boldsymbol{c}$ denotes the conditional input, typically a text embedding generated by a text encoder like CLIP (Radford et al. 2021) in text-to-image generation.
Following the training process, the diffusion model gradually generates $\boldsymbol{z}_{T-1}$, $\boldsymbol{z}_{T-2}$, · · ·, $\boldsymbol{z}_{0}$ from a randomly initialized noisy latent $\boldsymbol{z}_{T}$. Finally, the decoder of VAE is employed to generate the desired image $\boldsymbol{I}^{\prime}$ based on $\boldsymbol{z}_{0}$.
4.2 Cross-Modal Modeling Formulation
In our DiffX model, we firstly obtain the joint distribution by mapping the “RGB+X” modalities into the shared latent space. Specifically, during the forward process, the noising transition from $t-1$ to $t$ is conditioned on the joint modality $\mathbf{M}=\{\boldsymbol{m}_{1},\boldsymbol{m}_{2}\}$, which can be expressed as:
$q(\boldsymbol{z}_{t},\mathbf{M}|\boldsymbol{z}_{t-1})=q(\boldsymbol{z}_{t}|\boldsymbol{z}_{t-1},\mathbf{M})\boldsymbol{\Pi}_{i=1}^{2}q_{i}(\boldsymbol{m}_{i}).$ (2)
Meanwhile, the reverse process can be regarded as the reverse of the forward distributions, resulting in a joint distribution $p_{\boldsymbol{\theta}}(\boldsymbol{z}_{t-1},\mathbf{M}|\boldsymbol{z}_{t})$ at time step $t$:
$p_{\boldsymbol{\theta}}(\boldsymbol{z}_{t-1},\mathbf{M}|\boldsymbol{z}_{t})=p_{\boldsymbol{\theta}}(\boldsymbol{z}_{t-1}|\boldsymbol{z}_{t})\boldsymbol{\Pi}_{i=1}^{2}p_{\boldsymbol{\theta}}(\boldsymbol{m}_{i}|\boldsymbol{z}_{t}),$ (3)
where $\boldsymbol{\theta}$ denotes the parameters of the our DiffX model.
4.3 Dual-Path Variational AutoEncoder
As illustrated in Figure 4, the proposed Dual-Path Variational AutoEncoder (DP-VAE) is critical for the cross-modal generation task. In contrast to the conventional VAE used in LDM, our DP-VAE employs the form of “RGB+X” input and “RGB+X” output. In detail, the DP-VAE utilizes a single encoder $\mathcal{E}$ to encode the input $\boldsymbol{I}_{v},\boldsymbol{I}_{x}$} (denoted as $\{\boldsymbol{m}_{1},\boldsymbol{m}_{2}\}$ in Eq. (2)) into a latent representation $z$, and incorporates two parallel decoders $\mathcal{D}_{v}$ and $\mathcal{D}_{x}$ to generate the corresponding output $\boldsymbol{I}_{v}^{\prime},\boldsymbol{I}_{x}^{\prime}$}. The input images are processed through the separate convolutional layers, followed by element-wise addition before being fed to $\mathcal{E}$.
In addition, we have observed that image modality “X” usually exhibits distinct contour edges and prominent object positions. To enhance the reconstruction ability of the DP-VAE framework, we apply a Laplacian Pyramid (LP) (Burt and Adelson 1987) to extract the frequency-aware information from the input cross-modal image pairs. The extracted features are embedded into the various scale layers of the encoder through convolutional layers, enabling a comprehensive enhancement across various scales.
4.4 DiffX Model in Modality-Shared Space
The overall workflow of the DiffX model is shown in Figure 3. The noisy latent $\boldsymbol{z}_{T}$ is randomly initialized, which is adopted for generating $\boldsymbol{z}_{T-1}$, $\boldsymbol{z}_{T-2}$, · · ·, $\boldsymbol{z}_{0}$ by our DiffX-UNet. The diffusion processes involve grounding the effective conditional input, including the text captions and layout conditions. Finally, the parallel decoders $\mathcal{D}_{v}$ and $\mathcal{D}_{x}$ from the pre-trained DP-VAE are employed to generate the cross-modal “RGB+X” images based on $\boldsymbol{z}_{0}$, which is represented as the denoised joint distribution $p_{\boldsymbol{\theta}}(\boldsymbol{z}_{0},\mathbf{M}|\boldsymbol{z}_{1})$ in Eq. (3).
Figure 4: Illustration of our Dual-Path Variational AutoEncoder (DP-VAE). We adopt the Laplacian Pyramid (LP) strategy in LPTN (Liang, Zeng, and Zhang 2021).
Layout Condition Embedding.
Before the diffusion process, it is necessary to embed the layout conditions into grounding features to be fed into the DiffX-UNet.
1) For the box-based layout, we embed each box $\boldsymbol{b}_{i}=[x_{\text{min}},y_{\text{min}},x_{\text{max}},y_{\text{max}}]$ with label $\boldsymbol{l}_{i}$ into one grounding token. In detail, $\boldsymbol{b}_{i}$ and $\boldsymbol{l}_{i}$ are embedded via Fourier mapping $\mathcal{F}(\cdot)$ (Tancik et al. 2020) and the pre-trained CLIP text encoder $f_{\text{clip}}(\cdot)$, respectively. Afterwards, a Multi-Layer Perception (MLP) with parameters $\phi$ is adopted to encode the concatenated features:
$\boldsymbol{h}_{i}=\text{MLP}_{\phi}([\mathcal{F}(\boldsymbol{b}_{i}),f_{\text{clip}}(\boldsymbol{l}_{i})]);\boldsymbol{H}=[\boldsymbol{h}_{1},...,\boldsymbol{h}_{n}],$ (4)
where []delimited-[][$\cdot$][ ⋅ ] denotes the feature concatenation. The representation $\boldsymbol{H}$, which contains a total of $n$ embedded box features, is adopted as the layout conditional feature.
2) For the Seg., SOD, and edge layout $\boldsymbol{m}$, the pre-trained ConvNeXtψ model (Liu et al. 2022) is employed to extract the map features. After that, the Position Embedding $\boldsymbol{p}$ is added. Finally, an MLP with parameters $\varphi$ is used to generate the layout conditional feature:
$\boldsymbol{H}=\text{MLP}_{\varphi}(\mathcal{R}(\text{ConvNeXt}_{\psi}(\boldsymbol{m}))+\boldsymbol{p}),$ (5)
where $\mathcal{R}$ denotes the reshaping process to transform the feature into flattened tokens.
Long-CLIP Caption Embedding.
In our cross-modal generation task, the layout-based condition plays a crucial role in providing the general outline of objects within the images. However, text caption conditions $\boldsymbol{c}$ remain equally significant in cross-modal generation as they offer detailed descriptions of the images, encompassing aspects such as weather conditions, environments, surroundings, transportation modes, and more. Specifically, our task necessitates long text descriptions to capture the multitude of small objects and complexities present in the scene, as exemplified in Figure 2 (b). In our DiffX, we utilize the recent Long-CLIP model (Zhang et al. 2024) as the text encoder $f_{\text{L-clip}}(\cdot)$ to embed the text captions, which represents a significant advancement over its predecessor, CLIP, due to its ability to support long-text input. In detail, each caption $\boldsymbol{c}$ is embedded into a sequence of text embeddings $\boldsymbol{c}^{*}=f_{\text{L-clip}}(\boldsymbol{c})$.
Figure 5: Qualitative results of cross-modal “RGB+X” generation by our DiffX. The synthesis resolution is fixed to 512 $\times$ 512.
DiffX-UNet.
DiffX-UNet serves as the denoising model, featuring a combination of cascaded residual and Spatial-Transformer blocks. To enhance performance and efficiency, pre-trained weights from SD are employed for the frozen layers. The Spatial-Transformer block is illustrated in Figure 3, which is the fundamental component of the DiffX-UNet. It is composed of cascaded self-attention, cross-attention, and Feed Forward (FF) layers. The Feed Forward (FF) layer is an MLP with the middle dimensional expansion and skip connection. For the trainable FF layers in DiffX model, we add the scaling factors with the tanh activation. The cross-attention mechanism, derived from the self-attention mechanism, is formulated as follows:
$\text{Attn}(Q,K,V)=\text{softmax}\left(\frac{Q\cdot K^{T}}{\sqrt{d_{k}}}\right)\cdot V,$ (6)
where $Q$, $K$, and $V$ are derived from linear transformations of the features, and $d_{k}$ denotes the dimension of $K$. In our model, the visual tokens always serve as the modality to generate $Q$ in cross-attention layers.
In specific, the gated cross-attention layer and one FF layer are implemented prior to the Spatial-Transformer blocks, which aim to generate the enhanced layout conditional feature $\boldsymbol{H}^{*}$:
$\boldsymbol{H}^{\prime}=\boldsymbol{H}+\lambda\cdot\text{tanh}(\gamma_{1})\cdot\text{Cross-Attn}(\boldsymbol{H},\boldsymbol{c}^{*}),$ (7)
$\boldsymbol{H}^{*}=\boldsymbol{H}^{\prime}+\lambda\cdot\text{tanh}(\gamma_{2})\cdot\text{FF}(\boldsymbol{H}^{\prime}),$ (8)
where the hyperparameter $\lambda$ plays a crucial role in balancing quality and controllability. Additionally, $\gamma_{1}$ and $\gamma_{2}$ are two adaptive learnable scalars inside tanh activation.
Afterwards, the resulting $\boldsymbol{H}^{*}$ is fed into the gated self-attention within the Spatial-Transformer blocks. The gated self-attention, inspired by GLIGEN (Li et al. 2023), acts as an adapter module to effectively capture cross-modal relationships. Similarly, an FF is layer is integrated after each gated self-attention to build the long-range interaction:
$\boldsymbol{T}^{\prime}=\boldsymbol{T}+\mu\cdot\text{tanh}(\delta_{1})\cdot\mathcal{T}(\text{Self-Attn}([\boldsymbol{T},f_{\text{linear}}(\boldsymbol{H}^{*})])),$ (9)
$\boldsymbol{T}^{*}=\boldsymbol{T}^{\prime}+\mu\cdot\text{tanh}(\delta_{2})\cdot\text{FF}(\boldsymbol{T}^{\prime}),$ (10)
where $\mathcal{T}(\cdot)$ is a token selection operation that considers visual token positions of $\boldsymbol{T}$. The hyperparameter $\mu$ adjusts the weights of self-attention mechanism, whereas $\delta_{1}$ and $\delta_{2}$ are also two learnable scalars.
Compared with the Transformer blocks in SD, our DiffX-UNet incorporates Spatial-Transformer blocks that introduce an outer gated cross-attention layer and iterative gated self-attention layers. Moreover, we utilize the Long-CLIP caption embedding to enhance the layout features and the frozen cross-attention mechanism.
4.5 Training & Inference
Training. Algorithm 1 shows the optimization process.
1) The training of DP-VAE is critical prior to the training of DiffX for cross-modal generation, where we encode the cross-modal “RGB+X” image pair into the modality-shared latent space $\boldsymbol{z}$ and conduct end-to-end training. The training objective of DP-VAE includes the MSE loss and perceptual loss (Johnson, Alahi, and Fei-Fei 2016):
~$\tilde{\mathcal{L}}_{\text{DP}}\triangleq\mathcal{L}_{\text{mse}}(\mathbf{M},\widehat{\mathbf{M}})+\boldsymbol{\Sigma}_{i=1}^{2}\text{KL}\left(q_{i}(\boldsymbol{m}_{i})\|p_{\boldsymbol{\theta}}(\boldsymbol{m}_{i}|\boldsymbol{z})\right),$ (11)
where the second term denotes the perceptual loss fitted with Kullback-Leibler (KL) divergence.
2) The training objective of our DiffX model, namely the denoising function, can be formulated as:
~$\tilde{\mathcal{L}}_{\text{DiffX}}\triangleq\mathbb{E}_{q}[\boldsymbol{\Sigma}_{t>1}\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\boldsymbol{\vartheta}}(\boldsymbol{z}_{t},t,\boldsymbol{c}^{*},\boldsymbol{H}^{*})\right\|_{2}^{2}],$ (12)
where $t\sim\text{Uniform}(\{1,...,T\})$ is sampled in each iteration. Then, the latent $\boldsymbol{z}_{t}\sim q(\boldsymbol{z}_{t}|\boldsymbol{z}_{0},\mathbf{M})$ is sampled from the forward process with aggregated task information, which is fed into the reverse diffusion model to predict the forward noise.
Inference. The sampling steps are identical to those of the conventional LDM, with the only difference being the dual-branch decoding by the dual-path decoders of DP-VAE. The main purpose is to estimate cross-modal data with features from the corresponding heads of the reverse model.
Algorithm 1 Optimization process of DiffX.
0:  Training DP-VAE.
1:  While training:
2:      Prepare: “RGB+X” $\boldsymbol{I}_{v},\boldsymbol{I}_{x}$};
3:      Forward: $\{\boldsymbol{I}_{v}^{\prime},\boldsymbol{I}_{x}^{\prime}\}=\text{DP-VAE}(\{\boldsymbol{I}_{v},\boldsymbol{I}_{x}\})$;
4:      Minimize: ~$\tilde{\mathcal{L}}_{\text{DP}}$ formulated in Eq. (11);
5:      Update: DP-VAE;
6:  End While.
6:  Training DiffX (DiffX-UNet denoted as $\boldsymbol{\epsilon}_{\boldsymbol{\vartheta}}$).
7:  Define: $\alpha_{1}$, $\alpha_{2}$, …, $\alpha_{T}$ (derived from $\beta_{1}$, $\beta_{2}$, …, $\beta_{T}$);
8:  While training:
9:      Prepare: “RGB+X” $\boldsymbol{I}_{v},\boldsymbol{I}_{x}$}, caption $\boldsymbol{c}$, layout $\boldsymbol{b}/\boldsymbol{m}$;
10:      Forward: $\boldsymbol{z}_{0}=\text{DP-Encoder}(\{\boldsymbol{I}_{v},\boldsymbol{I}_{x}\})$;
11:      Forward: $\boldsymbol{c}^{*}$ and $\boldsymbol{H}^{*}$ based on $\boldsymbol{c}$ and $\boldsymbol{b}/\boldsymbol{m}$;
12:      Sample: $t\sim\text{Uniform}(\{1,...,T\})$, $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$;
13:      Calculate: $\boldsymbol{z}_{t}=\sqrt{\overline{\alpha}_{t}}\boldsymbol{z}_{0}+\sqrt{1-\overline{\alpha}_{t}}\boldsymbol{\epsilon}$;
14:      Minimize: $\mathbb{E}_{\boldsymbol{z}_{0},\boldsymbol{\epsilon},t}\left[\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\boldsymbol{\vartheta}}\left(\boldsymbol{z}_{t},t,\boldsymbol{c}^{*},\boldsymbol{H}^{*}\right)\|_{2}^{2}\right]$;
15:      Update: Trainable modules of DiffX;
16:  End While.
5 Experiments
5.1 RGB+X Datasets
We conduct experiments on three “RGB+X” datasets for four tasks to illustrate the generalizability of our DiffX. In detail, we adopt the clean-version FLIR dataset (F.A.Group 2019; Zhang et al. 2020) for Box $\rightarrow$ RGB+T and the MFNet dataset (Ha et al. 2017) for Segmentation (Seg.) maps $\rightarrow$ RGB+T. Meanwhile, the COME15K dataset (Zhang et al. 2021) is adopted for SOD $\rightarrow$ RGB+D and Edge $\rightarrow$ RGB+D.
5.2 Implementation Details
[…]
5.3 Experimental Results and Comparison
Unified Cross-Modal Generation.
As shown in Figure 5, we can see that DiffX is a unified framework for cross-modal generation under various layout conditions. The RGB+T generation benefits from its precise control over object positions and shapes. Meanwhile, it excels at RGB+D generation, generating highly realistic images with fine color and texture details. Most significantly, the generated X images exhibit strong alignment with the RGB images, demonstrating its ability to generate coherent pairs of cross-modal data.
Additionally, DiffX also enables the diverse cross-modal generation by editing the key captions, which is shown in Figure 6. In traffic-scene RGB+T generation, the generated weather condition can be changed by modifying keywords like “daytime” and “nighttime”, which is critical in autonomous driving systems. For the natural RGB+D generation, we can change the categories and attributes of the generated objects by adjusting the captions while keeping their structural integrity and outlines. It facilitates both creative manipulation and cross-modal image editing.
Figure 6: Diverse cross-modal generation by caption editing.
Table 1: Quantitative comparison with the baseline models.
[…]
Seg. $\rightarrow$ RGB + T VQVAE 0.673 0.588 91.58 98.61 0.437 0.596 0.508
VQGAN 0.552 0.489 70.24 81.05 0.518 0.659 0.630
DDPM 0.535 0.457 71.53 78.69 0.535 0.672 0.617
GLIGEN 0.548 0.428 69.77 73.58 0.520 0.667 0.641
DiffX 0.513 0.382 60.10 62.35 0.537 0.693 0.667
[…]
Edge $\rightarrow$ RGB + D VQVAE 0.411 0.285 46.82 55.80 0.721 0.167 0.653
VQGAN 0.389 0.290 40.82 41.58 0.733 0.159 0.747
DDPM 0.255 0.208 26.79 34.18 0.751 0.140 0.768
GLIGEN 0.249 0.211 25.07 34.06 0.743 0.152 0.762
DiffX 0.246 0.197 17.11 31.20 0.764 0.134 0.795
  • $\uparrow$/$\downarrow$ / ↓ denotes that higher / lower values are better.
Comparison with Existing Models.
To our knowledge, our DiffX is the first model for cross-modal “RGB+X” generation. Although the baseline models were designed for RGB image generation, we can modify them to have dual outputs for comparison. Considering the model adaptability, we compare the models in tasks of Seg. $\to$ RGB+T and Edge $\to$ RGB+D.
As shown in Figure 7, we can see that our DiffX outperforms the baseline methods significantly. The VQVAE (Van Den Oord, Vinyals et al. 2017) exhibits poor performance due to its unsuitability for dual-path outputs in end-to-end training. The VQGAN (Esser, Rombach, and Ommer 2021) and DDPM (Ho, Jain, and Abbeel 2020) can generate overall good images, but lack precision in capturing target objects and details. The modified GLIGEN (Li et al. 2023), which also performs diffusion training in a modality-shared latent space, yields superior results compared with other baseline models. However, it fails to generate the accurate background in the nighttime scene despite using captions for guidance. It proves the effectiveness of employing Long-CLIP for caption embedding and incorporating cross-attention mechanism in our DiffX model.
Figure 7: Qualitative comparison with the baseline models. The symbol denotes the modified versions.
On the other hand, we evaluate the generated image quality by LPIPS (Zhang et al. 2018) and FID (Heusel et al. 2017). Meanwhile, we adopt the pre-trained models in downstream tasks to evaluate the correspondence to the original layouts. We use EGFNet (Zhou et al. 2022) to test mean Intersection over Union (IoU) and mean Accuracy (Acc) on Seg. $\rightarrow$ RGB+T task, and use BBSNet (Fan et al. 2020) to test S-measure ($S_{\alpha}$) (Fan et al. 2017) and Mean Absolute Error ($M$) scores on Edge $\rightarrow$ RGB+D task. Additionally, we employ the SSIM score to evaluate the consistency of generated and ground-truth X images. The results in Table 1 demonstrate that our DiffX outperforms all baseline models in all metrics. Our impressive LPIPS and FID scores demonstrate the high-quality image generation by DiffX. Furthermore, the outstanding performance on downstream tasks highlights the effectiveness of DiffX in data augmentation for cross-modal vision tasks. Finally, the achieved SSIM scores on the X images (0.667 and 0.795 on the two tasks, respectively) demonstrate the advantage of our approach for consistent layout-aware generation.
Ablation Study.
Table 2: Ablation study on Laplacian Pyramid (LP) in DP-VAE for cross-modal image reconstruction.
($\uparrow$) Dataset Source DP-VAE w/o LP DP-VAE w/ LP
[…]
PSNR FLIR RGB+T 29.04 29.77 29.25 30.89
MFNet RGB+T 30.59 30.96 31.45 32.86
COME RGB+D 30.09 34.19 30.15 35.73
SSIM FLIR RGB+T 0.785 0.719 0.792 0.737
MFNet RGB+T 0.788 0.845 0.800 0.919
COME RGB+D 0.729 0.958 0.749 0.974
Table 3: Quantitative comparison with the umimodal DiffXs and cross-modal DiffX on generated image quality.
($\downarrow$) Task DiffXs DiffX
[…]
LPIPS Box $\rightarrow$ RGB+T 0.557 0.604 0.527 0.497
Seg. $\rightarrow$ RGB+T 0.551 0.424 0.513 0.382
SOD $\rightarrow$ RGB+D 0.718 0.489 0.579 0.305
Edge $\rightarrow$ RGB+D 0.258 0.215 0.246 0.197
FID Box $\rightarrow$ RGB+T 56.31 76.46 55.72 61.89
Seg. $\rightarrow$ RGB+T 71.58 75.17 60.10 62.35
SOD $\rightarrow$ RGB+D 71.23 80.54 51.59 52.83
Edge $\rightarrow$ RGB+D 19.36 35.60 17.11 31.20
Firstly, we conduct the ablation study on the Laplacian Pyramid (LP) in our DP-VAE. Results in Table 2 show that the LP structure effectively improves the PSNR and SSIM (Wang et al. 2004), facilitating cross-modal image reconstruction through frequency-aware incorporation.
Secondly, we also convert DiffX into a standard generative model (denoted as DiffXs) to generate RGB and X separately. We aim to compare its unimodal generation with the cross-modal RGB+X generation by DiffX. Quantitative comparison in Table 3 shows that the cross-modal results generated by DiffX exhibit better LPIPS and FID scores. This is because the diffusion training in modality-shared space can provide key complementary information for RGB and X image generation respectively.
Figure 8: Impact of text captions in cross-modal generation.
Thirdly, we conduct the ablation study on the impact of utilizing text captions on SOD $\rightarrow$ RGB+D and Seg. $\rightarrow$ RGB+T tasks, which heavily rely on text captions. Figure 8 shows that DiffX can effectively capture the crucial captions, while the model without caption embeddings generates misaligned images, even affecting the image quality. It also necessitates incorporating Long-Clip for embedding long text captions in cross-modal generation.
Adaptation to Multi-Modal Generation.
Given DiffX’s strong ability to generate cross-modal images, we also wonder if we can apply this framework to achieve more robust, controllable, and versatile generation across multiple modalities? Therefore, we conduct an experiment on COME15K dataset for “SOD $\rightarrow$ RGB+D+Edge” task to illustrate its adaptability. For this task, we modify the DP-VAE to support three-modality encoding and decoding, with the diffusion training carried out in a shared latent space that encompasses all three modalities.
Figure 9: Example results on “SOD $\rightarrow$ RGB+D+Edge” task.
The example results are shown in Figure 9. It is evident that the DiffX model can effectively accommodate the “SOD $\rightarrow$ RGB+D+Edge” task by generating multiple modalities. DiffX can accurately generate diverse generative results that align well with the SOD map and provided caption, producing corresponding RGB images based on the layout conditions and caption while also matching depth images and edge images effectively. Therefore, we are confident in the adaptability of DiffX for generating cohesive data across wide-range modalities. In the future, if any datasets with well-aligned “RGB+T+D” images occur, DiffX is a standout choice for layout-guided multi-modal generation.
6 Conclusion
In this study, we introduce the DiffX model for cross-modal “RGB+X”generation under various layout conditions. The key innovation of DiffX lies in its use of diffusion-based training within a shared latent space across modalities. Extensive experiments have demonstrated the robustness and versatility of our method, also showcasing its adaptability to multi-modal generation. Despite a slight increase in computational costs (less than 2%), our method maintains superior efficiency in terms of both time and storage compared to training separate models for individual modalities. Moving forward, our goal is to enhance DiffX to support multiple-to-multiple generative modeling, enabling the generation of “RGB+X+Y” or even more modalities. We aim to incorporate diverse layout conditions and facilitate simultaneous multi-modal generation based on user requirements.