### v2

# DiffX: Guide Your Layout to Cross-Modal Generative Modeling

Abstract
Diffusion models have made significant strides in language-driven and layout-driven image generation. However, most diffusion models are limited to visible RGB image generation. In fact, human perception of the world is enriched by diverse viewpoints, including chromatic contrast, thermal illumination, and depth information. In this paper, we introduce a novel diffusion model for general layout-guided cross-modal “RGB+X” generation, called DiffX. Firstly, we construct the cross-modal image datasets with text description by using LLaVA for image captioning, supplemented by manual corrections. Notably, DiffX presents a simple yet effective cross-modal generative modeling pipeline, which conducts diffusion and denoising processes in the modality-shared latent space, facilitated by our Dual-Path Variational AutoEncoder (DP-VAE). Moreover, we introduce the joint-modality embedder, which incorporates a gated cross-attention mechanism to link layout and text conditions. Meanwhile, the advanced Long-CLIP is employed for long caption embedding to improve user guidance. Through extensive experiments, DiffX demonstrates robustness and flexibility in cross-modal generation across three RGB+X datasets: FLIR, MFNet, and COME15K, guided by various layout types. It also shows the potential for adaptive generation of “RGB+X+Y” or more diverse modalities. Our code and constructed cross-modal image datasets are available at https://github.com/zeyuwang-zju/DiffX.
1 Introduction
Over the past decade, there have been significant advancements in image generation, driven by the improvement of deep generative models such as Variational AutoEncoders (VAEs) (Kingma and Welling 2013) and Generative Adversarial Networks (GANs) (Goodfellow et al. 2014). Notably, the recent diffusion models (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020) like DALL-E (Ramesh et al. 2021, 2022), Imagen (Saharia et al. 2022), and Stable Diffusion (SD) (Rombach et al. 2022) have gained popularity for their ability to generate high-fidelity images conditioned on text description or class labels. Moreover, the field has witnessed the rise of layout-to-image models based on various layout conditions, including bounding boxes, semantic maps, and keypoints (Xie et al. 2023; Li et al. 2023; Ma et al. 2024). The success of diffusion models can be attributed to their stable learning objective and the availability of large-scale multi-modal training data.
Figure 1: Cross-modal “RGB+X” generation in latent space.
Figure 2: (a) The process of constructing the image captions. (b) Examples of cross-modal images, labels, and prepared captions.
Current layout-to-image diffusion models primarily focus on generating high-quality visible RGB images. However, our perception of the world is enhanced by a multitude of perspectives that extend beyond the visible spectrum, including thermal imaging and depth perception. Cross-modal visual understanding typically leverages the input format of “RGB+X”, where X represents additional data like Thermal (T) or Depth (D) images. It offers a more comprehensive representation compared with RGB-only understanding, especially in challenging environments. At present, a significant limitation lies in the scarcity of cross-modal data containing pixel-aligned RGB+X image pairs. If the conventional generative models are adopted for cross-modal data augmentation, we can only generate RGB and X images separately, and as a result, the RGB and X images cannot form consistent pairs of matching data. Consequently, we wonder if we can generate the cross-modal images simultaneously under layout guidance by an integrated model?
In this work, we introduce a novel diffusion model for cross-modal image generation, called DiffX. As shown in Figure 1, our proposed DiffX model is capable of generating RGB+X image pairs guided by various layouts with user instructions. Our main contributions are as follows:
  • We propose an effective cross-modal generative modeling pipeline, which performs the diffusion and denoising processes in modality-shared latent space. In particular, the DP-VAE is implemented for cross-modal encoding.
  • We propose a joint-modality embedder to establish the comprehensive connection between layout and text conditions via the gated cross-attention mechanism. Besides, we employ the advanced Long-CLIP (Zhang et al. 2024) to capture critical information in cross-modal captions.
  • Due to the limited cross-modal data with text descriptions, we leverage a Large-Multimodal Model (LMM) to construct captions, followed by manual corrections.
  • Our experiments demonstrate that DiffX can generate high-quality and coherent “RGB+X” image pairs based on bounding boxes, semantic maps, salient maps, and edge contours. Additionally, it demonstrates adaptability to “RGB+X+Y” or diverse-modality generation.
Figure 3: The workflow of our DiffX model for cross-modal generation. It performs the diffusion and denoising processes in the modality-shared space. Finally, the denoised feature $\boldsymbol{z}_{0}$ is decoded into the “RGB+X” image pair via the dual-path decoders.
2 Related Works
2.1 Cross-Modal Visual Understanding
Cross-modal visual understanding normally adopts RGB+X input data. Thermal images have shown advantages in detecting objects in low-light conditions. For instance, CFT (Fang, Han, and Wang 2021) and ICAFusion (Shen et al. 2024) incorporate Transformer-based RGB+T feature fusion modules for multispectral object detection. In semantic segmentation, EGFNet (Zhou et al. 2022) leverages prior edge maps of RGB+T images to capture detailed contour information. Meanwhile, depth images exhibit valuable insights into spatial structures and 3D layouts. For example, HidaNet (Wu et al. 2023) utilizes a granularity-based attention mechanism to enhance the advantages of RGB+D features for Salient Object Detection (SOD). However, current cross-modal visual understanding is hindered by the lack of well-aligned RGB+X image pairs for training deep models.
2.2 Layout-to-Image Generation
Based on Denoising Diffusion Probabilistic Model (DDPM) (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020), diffusion models have undergone advancements in training and sampling techniques (Song, Meng, and Ermon 2021; Nichol and Dhariwal 2021). Recent studies have proposed layout-to-image diffusion models, which provide precise instructions for the object placement. Notably, LAW-Diffusion (Yang et al. 2023) integrates a spatial parser and adaptive guidance, enabling complex scene generation. BoxDiff (Xie et al. 2023) directly incorporates spatial conditions into the training-free denoising process. Additionally, InstanceDiffusion (Wang et al. 2024) focuses on precise instance-level control in image generation.
2.3 Multi-Modal Generation
Multi-modal generation has become a challenging task and attracted widespread research. MM-Diffusion (Ruan et al. 2023) integrates featuring two interconnected denoising autoencoders and a sequential multi-modal U-Net architecture for joint audio-video generation. Moreover, MM-Interleaved (Tian et al. 2024) dynamically extracts information from multi-scale features for image-text generation. In addition, ControlVAR (Li et al. 2024), MT-Diffusion (Chen et al. 2024), and 4M (Mizrahi et al. 2024) adopt large-scale natural datasets or private datasets for end-to-end training on conditional generation. Despite these advancements, there is currently no generative model specifically focused on layout-guided cross-modal “RGB+X” image generation.
3 Cross-Modal Dataset Construction
Since DiffX aims to sample cross-modal images under user instruction, it is crucial to construct datasets with accurate layout control and text description for model training. However, existing cross-modal datasets usually lack detailed text captions. In this section, we introduce the process of constructing the image captions, as illustrated in Figure 2 (a).
Caption Generation: To obtain high-quality image captions, we employ the advanced LMM, namely LLaVA-v1.5-7b model (Liu et al. 2024), to extract detailed descriptions of the images by the given prompt, including weather conditions, environmental settings, surrounding elements, transportation conditions, and identifiable objects.
Caption Correction: Despite the effectiveness of LLaVA in capturing visual content, it can sometimes generate inaccurate captions, especially in complex scenes. Therefore, in cases where the generated captions exhibit minimal errors, we conduct manual corrections to ensure the accuracy. However, instances of significant errors or ambiguous descriptions necessitate further corrections. Under such conditions, we utilize the GPT-4o model (Achiam et al. 2023) to rectify inaccurate information and provide additional descriptions.
By combining the capability of the LLaVA model with the manual correction and assistance of GPT-4o, the captioning process ensures fidelity and accuracy in describing the visual context in the cross-modal datasets. Examples of the data and prepared captions are shown in Figure 2 (b).
4 DiffX for Cross-Modal Generation
The overall workflow of the DiffX model is shown in Figure 3. The noisy latent $\boldsymbol{z}_{T}$ is randomly initialized, which is adopted for generating $\boldsymbol{z}_{T-1}$, $\boldsymbol{z}_{T-2}$, · · ·, $\boldsymbol{z}_{0}$ by our DiffX-UNet. The denoising process involves embedding the text captions and layout conditions for user instruction.
In the following subsections, we first introduce the preliminaries on Latent Diffusion Model (LDM). Then, we describe the workflow and components of our DiffX in detail.
4.1 Preliminaries on Latent Diffusion Model
Recently, LDM and its successor SD (Rombach et al. 2022) have pushed the boundaries of diffusion models by using a low-dimensional latent space. Inside LDM, a prior VAE is employed to extract the latent representation $\boldsymbol{z}$ of the ground-truth image $\boldsymbol{I}$. Subsequently, the basic denoising model $\boldsymbol{\epsilon}_{\boldsymbol{\vartheta}}$, normally implemented as a UNet (Ronneberger, Fischer, and Brox 2015) with residual (He et al. 2016) and self-attention (Vaswani et al. 2017) blocks, is adopted to obtain the noise $\boldsymbol{\epsilon}$ in the noisy latent $\boldsymbol{z}_{t}$ at time step $t$. The training objective can be represented as:
$\min_{\boldsymbol{\theta}}\mathcal{L}_{\text{LDM}}=\mathbb{E}_{\boldsymbol{z}_{t},\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t,\boldsymbol{c}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\boldsymbol{\vartheta}}\left(\boldsymbol{z}_{t},t,\boldsymbol{c}\right)\right\|_{2}^{2}\right],$ (1)
where $\boldsymbol{c}$ denotes the conditional input, typically a text embedding generated by a text encoder such as CLIP (Radford et al. 2021) in text-to-image generation.
Following the training process, the denoising model $\boldsymbol{\epsilon}_{\boldsymbol{\vartheta}}$ gradually generates $\boldsymbol{z}_{T-1}$, $\boldsymbol{z}_{T-2}$, · · ·, $\boldsymbol{z}_{0}$ from a randomly initialized noisy latent $\boldsymbol{z}_{T}$. Finally, the decoder of VAE is employed to generate the target image $\boldsymbol{I}^{\prime}$ based on $\boldsymbol{z}_{0}$.
4.2 Cross-Modal Modeling Formulation
In our DiffX model, we first obtain the joint-modality distribution by mapping the “RGB+X” images into the shared latent space. Specifically, during the forward process, the noising transition from $t-1$ to $t$ is conditioned on the joint modality $\mathbf{M}=\{\boldsymbol{m}_{1},\boldsymbol{m}_{2}\}$, which can be expressed as:
$q(\boldsymbol{z}_{t},\mathbf{M}|\boldsymbol{z}_{t-1})=q(\boldsymbol{z}_{t}|\boldsymbol{z}_{t-1},\mathbf{M})\boldsymbol{\Pi}_{i=1}^{2}q_{i}(\boldsymbol{m}_{i}).$ (2)
Meanwhile, the reverse process can be regarded as the converse of the forward distributions, resulting in a joint distribution $p_{\boldsymbol{\theta}}(\boldsymbol{z}_{t-1},\mathbf{M}|\boldsymbol{z}_{t})$ at time step $t$:
$p_{\boldsymbol{\theta}}(\boldsymbol{z}_{t-1},\mathbf{M}|\boldsymbol{z}_{t})=p_{\boldsymbol{\theta}}(\boldsymbol{z}_{t-1}|\boldsymbol{z}_{t})\boldsymbol{\Pi}_{i=1}^{2}p_{\boldsymbol{\theta}}(\boldsymbol{m}_{i}|\boldsymbol{z}_{t}),$ (3)
where $\boldsymbol{\theta}$ denotes the parameters of the DiffX model.
4.3 Dual-Path Variational AutoEncoder
As illustrated in Figure 4, the proposed DP-VAE is critical for our cross-modal generative task. In contrast to the conventional VAE used in LDM, our DP-VAE employs the input form of “RGB+X” and output form of “RGB+X”. In detail, the DP-VAE utilizes a single encoder $\mathcal{E}$ to encode the input $\boldsymbol{I}_{v},\boldsymbol{I}_{x}$} (denoted as $\{\boldsymbol{m}_{1},\boldsymbol{m}_{2}\}$ in Eq. (2)) into the modality-shared latent representation $z$. The input images are processed through the separate convolutional layers, followed by element-wise addition before being fed to $\mathcal{E}$. Then, it incorporates two parallel decoders $\mathcal{D}_{v}$ and $\mathcal{D}_{x}$ to generate the corresponding output $\boldsymbol{I}_{v}^{\prime},\boldsymbol{I}_{x}^{\prime}$}.
In addition, we have observed that image modality “X” usually exhibits distinct contour edges and prominent object positions. To enhance the reconstruction ability of the DP-VAE, we apply a Laplacian Pyramid (LP) (Burt and Adelson 1987) to extract the high-frequency information from the input cross-modal image pairs. Subsequently, the extracted features are embedded into the multi-scale layers of the encoder through cascaded convolutional layers, enabling a comprehensive enhancement at various frequencies.
During the inference phase of DiffX, the parallel decoders $\mathcal{D}_{v}$ and $\mathcal{D}_{x}$ from pre-trained DP-VAE are employed to generate the cross-modal “RGB+X” images based on the denoised latent feature $\boldsymbol{z}_{0}$.
Figure 4: Illustration of our Dual-Path Variational AutoEncoder (DP-VAE). We adopt the Laplacian Pyramid (LP) strategy in LPTN (Liang, Zeng, and Zhang 2021).
4.4 Joint-Modality Embedder
Before the denoising process, it is necessary to embed the layout and text conditions into grounding features. To establish a comprehensive connection between layout and text conditions, we propose the joint-modality embedder ($\mathcal{J}_{\omega}$):
$\boldsymbol{H}^{*},\boldsymbol{c}^{*}=\mathcal{J}_{\omega}(\text{Layout},\text{Text}).$ (4)
where $\boldsymbol{H}^{*}$ is the text-aware layout feature, and $\boldsymbol{c}^{*}$ is the embedded caption feature. In detail, $\mathcal{J}_{\omega}$ is composed of the layout encoder, Long-CLIP, a gated cross-attention layer, and a Feed Forward (FF) layer. The FF layer is an MLP with middle-dimensional expansion.
Layout Condition Embedding.
1) For the box-based layout, boxes $\boldsymbol{b}_{i}$ and labels $\boldsymbol{l}_{i}$ are embedded via Fourier mapping $\mathcal{F}(\cdot)$ (Tancik et al. 2020) and CLIP text encoder $f_{\text{clip}}(\cdot)$, respectively. Then, a Multi-Layer Perception (MLP) with parameters $\phi$ is used to encode them into grounding tokens:
$\boldsymbol{h}_{i}=\text{MLP}_{\phi}([\mathcal{F}(\boldsymbol{b}_{i}),f_{\text{clip}}(\boldsymbol{l}_{i})]);\boldsymbol{H}=[\boldsymbol{h}_{1},...,\boldsymbol{h}_{n}],$ (5)
where []delimited-[][$\cdot$][ ⋅ ] denotes the feature concatenation. The representation $\boldsymbol{H}$, which contains a total of $n$ embedded box features, is adopted as the layout conditional feature.
2) For the semantic mask layout represented by $\boldsymbol{m}$, we utilize the the pre-trained ConvNeXtψ model (Liu et al. 2022) to extract the in-depth semantic feature. Subsequently, the position embedding $\boldsymbol{p}$ is added, and an MLP with parameters $\varphi$ is used to generate the layout feature:
$\boldsymbol{H}=\text{MLP}_{\varphi}(\mathcal{R}(\text{ConvNeXt}_{\psi}(\boldsymbol{m}))+\boldsymbol{p}),$ (6)
where $\mathcal{R}$ denotes the process of reshaping and flattening the semantic feature into tokens.
Long-CLIP Caption Embedding.
Text conditions $\boldsymbol{c}$ remain equally significant in cross-modal generation as they provide detailed image descriptions, including various aspects such as weather, environments, and transportation. Specifically, our task necessitates long text captions to describe the small objects and complex backgrounds present in the cross-modal scene, as shown in Figure 2 (b). In our DiffX, we utilize the advanced Long-CLIP as the text encoder, which represents notable improvement over its predecessor, CLIP, due to its ability to support text inputs of up to 248 tokens. In detail, each caption $\boldsymbol{c}$ is embedded into a sequence of text embeddings $\boldsymbol{c}^{*}=f_{\text{L-clip}}(\boldsymbol{c})$.
Gated Cross-Attention.
Inside our $\mathcal{J}_{\omega}$, the gated Cross-Attention (CA) layer is implemented to fuse the layout and text conditional features, which is formulated as follows:
$\text{CA}(\boldsymbol{H},\boldsymbol{c}^{*})=\text{softmax}\left(\frac{W_{q}(\boldsymbol{H})\cdot W_{k}(\boldsymbol{c}^{*})^{T}}{\sqrt{d_{k}}}\right)\cdot W_{v}(\boldsymbol{c}^{*}),$ (7)
where $W_{q}$, $W_{k}$, and $W_{v}$ denote the linear transformations, and $d_{k}$ denotes the dimension of $W_{k}(\boldsymbol{c}^{*})$. Here, the gated cross-attention derives the query from layout condition $\boldsymbol{H}$, while deriving key and value from text condition $\boldsymbol{c}^{*}$.
In detail, a gated CA layer and an FF layer are sequentially implemented in $\mathcal{J}_{\omega}$, which aim to generate the text-aware layout conditional feature $\boldsymbol{H}^{*}$:
$\boldsymbol{H}^{\prime}=\boldsymbol{H}+\lambda\cdot\text{tanh}(\gamma_{1})\cdot\text{CA}(\boldsymbol{H},\boldsymbol{c}^{*}),$ (8)
$\boldsymbol{H}^{*}=\boldsymbol{H}^{\prime}+\lambda\cdot\text{tanh}(\gamma_{2})\cdot\text{FF}(\boldsymbol{H}^{\prime}),$ (9)
where the hyperparameter $\lambda$ plays a crucial role in balancing feature quality and controllability. Additionally, $\gamma_{1}$ and $\gamma_{2}$ are two adaptive learnable scalars inside the tanh activation.
Figure 5: Qualitative results of cross-modal “RGB+X” generation by our DiffX. The synthesis resolution is fixed to 512 $\times$ 512.
4.5 DiffX-UNet
As illustrated in Figure 3, DiffX-UNet serves as the basic model for the denoising process in the modality-shared latent space. It is composed of cascaded residual blocks and Spatial-Transformer blocks. The Spatial-Transformer block is the critical component of the DiffX-UNet, consisting of cascaded attention-based layers and FF layers. To enhance the training efficiency, pre-trained model weights from SD-v1-4 are initialized for the frozen layers.
The joint-modality feature $\boldsymbol{H}^{*}$ embedded by $\mathcal{J}_{\omega}$ is fed into the gated Self-Attention (SA) layer within the Spatial-Transformer blocks, which acts as an adapter module to capture the cross-modal relationships. Similar to the gated CA layer, an FF layer is integrated after each gated SA layer to build the long-range interaction:
$\boldsymbol{T}^{\prime}=\boldsymbol{T}+\mu\cdot\text{tanh}(\delta_{1})\cdot\mathcal{T}(\text{SA}([\boldsymbol{T},f_{\text{linear}}(\boldsymbol{H}^{*})])),$ (10)
$\boldsymbol{T}^{*}=\boldsymbol{T}^{\prime}+\mu\cdot\text{tanh}(\delta_{2})\cdot\text{FF}(\boldsymbol{T}^{\prime}),$ (11)
where $\mathcal{T}(\cdot)$ is a token selection operation that selects the visual token positions of $\boldsymbol{T}$. The hyperparameter $\mu$ adjusts the weights of the SA mechanism, whereas $\delta_{1}$ and $\delta_{2}$ are also two learnable scalars. For the trainable FF layers, we also add the scaling factors with the tanh activation.
Compared with the conventional Transformer blocks in SD, our Spatial-Transformer blocks integrate the iterative gated SA layers to adaptively embed the joint-modality feature $\boldsymbol{H}^{*}$. Moreover, the text condition $\boldsymbol{c}^{*}$ embedded by Long-CLIP further enhances the long caption awareness.
4.6 Training & Inference
Training.
Algorithm 1 shows the optimization process.
1) The training of DP-VAE is critical prior to the training of DiffX for cross-modal generation, where we encode the cross-modal “RGB+X” image pair into the modality-shared latent space $\boldsymbol{z}$ and conduct end-to-end reconstruction. The training objective of DP-VAE includes the MSE loss and perceptual loss (Johnson, Alahi, and Fei-Fei 2016):
~$\tilde{\mathcal{L}}_{\text{DP}}\triangleq\mathcal{L}_{\text{mse}}(\mathbf{M},\widehat{\mathbf{M}})+\boldsymbol{\Sigma}_{i=1}^{2}\text{KL}\left(q_{i}(\boldsymbol{m}_{i})\|p_{\boldsymbol{\theta}}(\boldsymbol{m}_{i}|\boldsymbol{z})\right),$ (12)
where the second term denotes the perceptual loss fitted with Kullback-Leibler (KL) divergence.
2) The training objective of our DiffX model, namely the latent denoising function, can be formulated as:
~$\tilde{\mathcal{L}}_{\text{DiffX}}\triangleq\mathbb{E}_{q}[\boldsymbol{\Sigma}_{t>1}\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\boldsymbol{\vartheta}}(\boldsymbol{z}_{t},t,\boldsymbol{c}^{*},\boldsymbol{H}^{*})\right\|_{2}^{2}],$ (13)
where $t\sim\text{Uniform}(\{1,...,T\})$ is sampled in each iteration. Then, the latent $\boldsymbol{z}_{t}\sim q(\boldsymbol{z}_{t}|\boldsymbol{z}_{0},\mathbf{M})$ is initialized, which is fed into the DiffX model to predict the forward noise.
Inference.
The difference in the sampling steps between our DiffX and conventional LDM is the dual-branch decoding by our DP-VAE. The main purpose is to estimate cross-modal images from features of the corresponding latent distributions of the DiffX model.
Algorithm 1 Optimization process of DiffX.
0:  Training DP-VAE.
1:  While training:
2:      Prepare: “RGB+X” $\boldsymbol{I}_{v},\boldsymbol{I}_{x}$};
3:      Forward: $\{\boldsymbol{I}_{v}^{\prime},\boldsymbol{I}_{x}^{\prime}\}=\text{DP-VAE}(\{\boldsymbol{I}_{v},\boldsymbol{I}_{x}\})$;
4:      Minimize: ~$\tilde{\mathcal{L}}_{\text{DP}}$ formulated in Eq. (12);
5:      Update: DP-VAE;
6:  End While.
6:  Training DiffX (DiffX-UNet denoted as $\boldsymbol{\epsilon}_{\boldsymbol{\vartheta}}$).
7:  Define: $\alpha_{1}$, $\alpha_{2}$, …, $\alpha_{T}$ (derived from $\beta_{1}$, $\beta_{2}$, …, $\beta_{T}$);
8:  While training:
9:      Prepare: “RGB+X” $\boldsymbol{I}_{v},\boldsymbol{I}_{x}$}, caption $\boldsymbol{c}$, layout $\boldsymbol{b}/\boldsymbol{m}$;
10:      Forward: $\boldsymbol{z}_{0}=\text{DP-Encoder}(\{\boldsymbol{I}_{v},\boldsymbol{I}_{x}\})$;
11:      Forward: $\boldsymbol{c}^{*}$ and $\boldsymbol{H}^{*}$ based on $\boldsymbol{c}$ and $\boldsymbol{b}/\boldsymbol{m}$;
12:      Sample: $t\sim\text{Uniform}(\{1,...,T\})$, $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$;
13:      Calculate: $\boldsymbol{z}_{t}=\sqrt{\overline{\alpha}_{t}}\boldsymbol{z}_{0}+\sqrt{1-\overline{\alpha}_{t}}\boldsymbol{\epsilon}$;
14:      Minimize: $\mathbb{E}_{\boldsymbol{z}_{0},\boldsymbol{\epsilon},t}\left[\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\boldsymbol{\vartheta}}\left(\boldsymbol{z}_{t},t,\boldsymbol{c}^{*},\boldsymbol{H}^{*}\right)\|_{2}^{2}\right]$;
15:      Update: Trainable modules of DiffX;
16:  End While.
5 Experiments
5.1 Cross-Modal Datasets
We conduct experiments on three “RGB+X” datasets for five cross-modal generative tasks. In detail, we adopt the clean-version FLIR dataset (F.A.Group 2019; Zhang et al. 2020) for Box $\rightarrow$ RGB+T and the MFNet dataset (Ha et al. 2017) for Segmentation (Seg.) maps $\rightarrow$ RGB+T. Meanwhile, the COME15K dataset (Zhang et al. 2021) is adopted for SOD $\rightarrow$ RGB+D, Edge $\rightarrow$ RGB+D, and SOD $\rightarrow$ RGB+D+Edge.
5.2 Implementation Details
[…]
5.3 Experimental Results and Comparison
Unified Cross-Modal Generation.
As shown in Figure 5, we can see that DiffX is a unified framework for cross-modal generation under various layout conditions. The RGB+T generation benefits from its precise control over object positions and shapes. Meanwhile, it excels at RGB+D generation, generating highly realistic images with fine color and texture details. Most significantly, the generated X images exhibit strong alignment with the RGB images, demonstrating its ability to generate coherent pairs of cross-modal data.
Additionally, DiffX also enables the diverse cross-modal generation by editing the key captions, as shown in Figure 6. In traffic-scene RGB+T generation, the generated weather condition can be changed by modifying keywords like “daytime” and “nighttime”, which is critical in autonomous driving systems. For the natural RGB+D generation, we can change the categories and attributes of the generated objects by adjusting the captions while keeping their structural integrity and outlines. Therefore, it facilitates both creative manipulation and cross-modal image editing.
Figure 6: Diverse cross-modal generation by caption editing.
Comparison with Existing Models.
To our knowledge, our DiffX is the first model for cross-modal “RGB+X” generation. Although the baseline models were designed for RGB image generation, we can modify them to have dual outputs for comparison. Considering the model adaptability, we compare the models in tasks of Seg. $\to$ RGB+T and Edge $\to$ RGB+D.
As shown in Figure 7, we can see that our DiffX significantly outperforms the baseline methods in qualitative results. The VQVAE (Van Den Oord, Vinyals et al. 2017) exhibits poor performance due to its unsuitability for dual-path outputs in end-to-end training. The VQGAN (Esser, Rombach, and Ommer 2021) and DDPM (Ho, Jain, and Abbeel 2020) can generate overall good images but lack precision in capturing target objects and details. The modified GLIGEN (Li et al. 2023), which also performs diffusion training in a modality-shared latent space, yields superior results compared with other baseline models. However, it fails to generate accurate backgrounds in the nighttime scene despite using captions for guidance. It proves the effectiveness of employing Long-CLIP for caption embedding and joint-modality embedder for modal connection.
Figure 7: Qualitative comparison with the baseline models. The symbol denotes the modified versions.
Table 1: Quantitative comparison with the baseline models.
[…]
Seg. $\rightarrow$ RGB + T VQVAE 0.673 0.588 91.58 98.61 0.437 0.596 0.508
VQGAN 0.552 0.489 70.24 81.05 0.518 0.659 0.630
DDPM 0.535 0.457 71.53 78.69 0.535 0.672 0.617
GLIGEN 0.548 0.428 69.77 73.58 0.520 0.667 0.641
DiffX 0.513 0.382 60.10 62.35 0.537 0.693 0.667
[…]
Edge $\rightarrow$ RGB + D VQVAE 0.411 0.285 46.82 55.80 0.721 0.167 0.653
VQGAN 0.389 0.290 40.82 41.58 0.733 0.159 0.747
DDPM 0.255 0.208 26.79 34.18 0.751 0.140 0.768
GLIGEN 0.249 0.211 25.07 34.06 0.743 0.152 0.762
DiffX 0.246 0.197 17.11 31.20 0.764 0.134 0.795
  • $\uparrow$/$\downarrow$ / ↓ denotes that higher / lower values are better.
On the other hand, we evaluate the generated image quality by LPIPS (Zhang et al. 2018) and FID (Heusel et al. 2017). Meanwhile, we adopt the pre-trained models in downstream tasks to evaluate the correspondence to the original layouts. We use EGFNet (Zhou et al. 2022) to test mean Intersection over Union (IoU) and mean Accuracy (Acc) on Seg. $\rightarrow$ RGB+T task, and use BBSNet (Fan et al. 2020) to test S-measure ($S_{\alpha}$) (Fan et al. 2017) and Mean Absolute Error ($M$) scores on Edge $\rightarrow$ RGB+D task. Additionally, we employ the SSIM score to evaluate the consistency of generated and ground-truth X images. The results in Table 1 demonstrate that our DiffX outperforms all baseline models in all metrics. Our impressive LPIPS and FID scores demonstrate the high-quality image generation by DiffX. Furthermore, the outstanding performance on downstream tasks highlights the effectiveness of DiffX in data augmentation for cross-modal vision tasks. Finally, the achieved SSIM scores on the X images (0.667 and 0.795 on the two tasks, respectively) illustrate the advantage of our approach for consistent layout-aware cross-modal generation.
Ablation Study.
Table 2: Ablation study on Laplacian Pyramid (LP) in DP-VAE for cross-modal image reconstruction.
($\uparrow$) Dataset Source DP-VAE w/o LP DP-VAE w/ LP
[…]
PSNR FLIR RGB+T 29.04 29.77 29.25 30.89
MFNet RGB+T 30.59 30.96 31.45 32.86
COME RGB+D 30.09 34.19 30.15 35.73
SSIM FLIR RGB+T 0.785 0.719 0.792 0.737
MFNet RGB+T 0.788 0.845 0.800 0.919
COME RGB+D 0.729 0.958 0.749 0.974
Table 3: Quantitative comparison with the umimodal DiffXs and cross-modal DiffX on generated image quality.
($\downarrow$) Task DiffXs DiffX
[…]
LPIPS Box $\rightarrow$ RGB+T 0.557 0.604 0.527 0.497
Seg. $\rightarrow$ RGB+T 0.551 0.424 0.513 0.382
SOD $\rightarrow$ RGB+D 0.718 0.489 0.579 0.305
Edge $\rightarrow$ RGB+D 0.258 0.215 0.246 0.197
FID Box $\rightarrow$ RGB+T 56.31 76.46 55.72 61.89
Seg. $\rightarrow$ RGB+T 71.58 75.17 60.10 62.35
SOD $\rightarrow$ RGB+D 71.23 80.54 51.59 52.83
Edge $\rightarrow$ RGB+D 19.36 35.60 17.11 31.20
Firstly, we conduct the ablation study on the Laplacian Pyramid (LP) in our DP-VAE. Results in Table 2 show that the LP structure effectively improves the PSNR and SSIM (Wang et al. 2004), facilitating cross-modal reconstruction through frequency information incorporation.
Secondly, we aim to compare the unimodal generation with the cross-modal RGB+X generation by DiffX. Therefore, we convert DiffX into a standard generative model (denoted as DiffXs) to generate RGB and X separately. Quantitative comparison in Table 3 shows that the cross-modal results generated by DiffX exhibit better LPIPS and FID scores. This is because the diffusion and denoising training in modality-shared space can provide key complementary information for RGB and X image generation, respectively.
Figure 8: Impact of text captions in cross-modal generation.
Thirdly, we conduct the ablation study on the impact of text captions on SOD $\rightarrow$ RGB+D and Seg. $\rightarrow$ RGB+T tasks. The qualitative comparison in Figure 8 shows that DiffX can effectively capture the crucial captions, while the variant model without caption embeddings generates broken or misaligned images, ultimately affecting the image quality. Therefore, it also necessitates incorporating Long-CLIP for embedding long text captions in cross-modal generation.
Adaptation to Multi-Modal Generation.
Given DiffX’s strong ability to generate cross-modal images, we also wonder if we can apply this framework to achieve more robust, controllable, and versatile generation across multiple modalities? Therefore, we conduct an experiment on COME15K dataset for “SOD $\rightarrow$ RGB+D+Edge” task to illustrate its adaptability. For this task, we modify the DP-VAE to support three-modality encoding and decoding, with the denoising training carried out in a shared latent space that encompasses three modalities.
Figure 9: Example results on “SOD $\rightarrow$ RGB+D+Edge” task.
The example qualitative results are shown in Figure 9. It is evident that the DiffX model can effectively adapt to the “SOD $\rightarrow$ RGB+D+Edge” task. DiffX can accurately generate diverse modal images that align well with the provided SOD maps and captions, producing corresponding RGB images based on the layout conditions while also matching depth images and edge images effectively. Therefore, DiffX is a standout choice for generating cohesive data across wide-range modalities. We also believe it can extend to cross-modal video and 3D image generative modeling in modality-shared latent diffusion.
6 Conclusion
In this study, we introduce the DiffX model for cross-modal “RGB+X”generation under various layout conditions. The key innovation of DiffX lies in its use of diffusion-based training within a shared latent space across modalities. Extensive experiments have demonstrated the robustness and versatility of our method, also showcasing its adaptability to multi-modal generation. Despite a slight increase in computational costs (less than 1%), our method maintains superior efficiency and storage when compared to training separate models for generating individual modalities. Moving forward, our goal is to enhance DiffX to support multiple-to-multiple generative modeling, enabling the generation of “RGB+X+Y” or even more modalities. We also aim to incorporate diverse layout conditions and facilitate simultaneous multi-modal generation based on user instruction.